OpenStudy (anonymous):

1) Solve: 1. 6y + 2 – 4y = -18 2. 8(t + 7) = 32

OpenStudy (anonymous):

2y+2=-18 2y=-20 y=-10 8t+8*7=24 8t+56=24 8t=-32 t=-4